Fiber Optics

Thursday, March 1st, 2007

DSC00076Fiber Creations of Arizona came out today and glued the fiber optics back into the waterfall, again. This time he said they would also silicone around it to keep it from falling out, again. He knew nothing about having to replace it as was discussed in previous meetings.

Replaster Started

Tuesday, February 27th, 2007

DSC00068Prado and Sons came out and started chipping away at the plaster. They are removing the plaster from around all the fixtures and along the top edge. They will then apply a adhesive layer and then a finish coat.
I called Debbie Moles and was wondering why they were going to refinish the pool before repairing the waterfall and fiber optics.
She called me back and said that she called Prado and had them wait to finish till all the other work has been completed.
The fiber optics should be installed on Thursday or Friday. Allen tile will be out Friday at 8 am to look over replacing the stucco with tile.

Repairs

Wednesday, January 24th, 2007

Kelly Sharp with Paddock Pools showed up along with a diver to fix the main drain and cleaning jets. He also replaced the filter air bleed and pressure gage due to it was allowing air back into the system. They also blew out a chunk of plaster from the main drain, it is no longer plugged. The gate to the pool area was adjusted so it closes now.
Dave with Aqua Torque came out and looked at the footprint etched into the step. He tried blending it with the surrounding area, but was ineffective. He said they might be able to remove it by draining and burning it off with acid. But he could not guarantee it would work. The other areas in the plaster that are stained or rough can be fixed.
Fiberstar looked at the fiber optics falling out of the waterfall. He would have to get in contact with the installers to see if they had any stronger glue to hold it in. If not they may have to redo the optic in the water feature.
I asked Kelly Sharp if he knew when they were going to fix the stucco on the face of the waterfall. He did not know. I mentioned that the pool heater would not heat and that it freezes over even when it is warm outside. I also told him that all work has to be completed before this years swimming season. He would not commit to a time frame. So I am going to make my own. If the pool repairs are not complete before March 31, 2007 I am turning everything over to lawyers to battle it out. Unfortunately as soon as lawsuits are filed all work stops till settled but not much work has been done anyway.

Another Pump

Tuesday, June 13th, 2006

The service tech came out and replaced the wok pot pump. He also blew out the main drain again, removing several large chunks of plaster from inside the line.
The fiber optics in the waterfall was glued back in place again.

Call

Monday, June 12th, 2006

DSC00004Kimberly called and said she had been told that the fiber optics had been repaired. We went outside and looked, but the fiber optics was still hanging out of the waterfall. She will call us back when she finds out what happened to the fiber optic people.

She called back and said they will be out within a few days to take care of the fiber optics.
I asked her about the tile people because the deck is lifting due to water seeping between the tile and deck. She will look into it and let me know when they will be out.

Call

Friday, June 2nd, 2006

Called again and left another message for Debbie.
Called back and talked with Rebecca. She will have someone out Tuesday to look at the main drain and air in system. The tile company will call to set up a time for caulking. The fiber optic people will call to set a time for the fiber optics to come out.
Electritions called and had the wrong work order. They wanted to put in controls for the lights. I told them that had been done a long time ago. Told them about the wok pot pump going on and off, and they said it sounded like the pump was failing. I guess they don’t do pumps.

Today’s list

Tuesday, May 30th, 2006

Called Paddock Pools warreanty dept. Rebecca transferred me to Debbie Moles voice mail. Waiting for call back.
Today’s list:
Wok pot pump shutting off and resetting
Fiber optic lines falling out of waterfall
Caulk between deck and waterline tile separating again
Stains still there
Main drain still pulling from side port and not much from bottom drain
Still sucking air into main filter pump, even when ozone shut off
Deck lifting and crack at deep end skimmer

Falling

Monday, May 29th, 2006

Fiber optic lines falling out of waterfall again. Caulk between deck and waterline tile seperating also.

Fiber Optics

Monday, May 1st, 2006

DSC00304The fiber optic people came out and glued the fiber optics back into the waterfall.
I changed out the rock in the wok pots from river rock to a smaller beach pebble. It now matches the look of the pots in Paddocks showroom wok pots. I cut down the pipes in the wok pots below the water line. It now flows more water while keeping a bubbling type effect.

Repairs

Friday, April 28th, 2006

Electrician came out and set the fiber optics on the remote. Also seperated the color wheels.
Plastic pipe scrap removed from under wok pots.
Another Jandy valve top was replaced.

Diving Board

Saturday, April 15th, 2006

DSC00206They set the diving board and filled in the wok pot fiber optic cut.

Clog

Tuesday, March 28th, 2006

Cody cleaned out the main drain. Plaster and junk was stuck down in the pipe. He said there was still some suction from the side safety drain but that may be normal.
A company looked at the crooked diving board. He said it did not look like they would have to tear out the deck to straighten it.
I talked with Fiberstars about getting the lights to sync. They replied that with the X-10 remote you can have synced or independent but not both. I guess I will have to live with independent lighting for now. So this narrows the list down by eliminating # 2, # 4, # 6.
Spoke with fiber optic people about the fiber on wok pot # 3 being in conduit. With the bends at 90 degrees, they would not be able to pull them out if they needed to repair them anyways. So the verdict for that one is it doesn’t matter if it is in conduit, future repairs will just mean rebuilding wok pot bases.

Glue Job

Monday, March 27th, 2006

DSC00159Fiber optics under waterfall glued in. Looks like a glue job. I will have to rebuild the top of the wall so it has a cantilever edge like the rest of the pool to hide the sheer. I don’t trust Poco Verde to try and fix it. The fiber optic controller is not only the wrong one but was not installed per manufacturer’s instructions. It was placed below the water line of the wok pots, possibly allowing the water to flow from the wok pots back into the controller, shorting it out. He said that he would let the office know about the plugged wok pot bubblers. Also they need to have the electrician come out and wire the controller into the remote. The Jazz lights no longer sync.
DSC00164Person came by to fix the pool stains. Stains, also known as perma-dirt, are still visible. Rust spot near waterfall, several streaks down the sides of the pool where the water line tile bled down. Plaster has dirt embedded in it throughout the surface. They plastered it right before a dust storm.
Cody called and said he could not make it today, still no air for his tank. The main drain remains plugged.
I found where the stucco on the waterfall is separating from the wall. A crack has formed where the tile meets the stucco and is lifting away from the wall on both sides.
This week’s list:
1. Stucco lifting from waterfall wall.
2. Lighting color wheels won’t sync.
3. Color wheels not able to change separately.
4. Pool lighting will not sync with fiber optic lighting.
5. Fiber optic controller not on remote.
6. Fiber optic controller not an auto-sync.
7. Stains in plaster.
8. Stains and chips on deck.
9. Grout on water level tile.
10. No conduit in concrete riser for wok pot # 3 fiber optics.
11. Wok pot lines plugged and do not spray correctly.
12. Diving board set crooked.
13. Gauges for cleaning system rusted.
14. Gauges for pumps broken.
15. Main drain plugged.
16. Water leak at Jandy valve for one of the wok pots.
17. U/V, Salt controller, timer, mounting poles falling down.

Lights

Friday, March 24th, 2006

DSC00142The electrician was out and wired the pool lights. He replaced the F/X light with Jazz lights. They can now be turned off and on separately, but the colors have to be the same.
Cody was out and fixed the cleaning system, replaced two gauges, and tried to clear out the main drain. He will be back on Monday to pull the main drain and clean out the line that is plugged.
DSC00144Scott Dilliam came out to look over the progress. I showed him where the fiber optics is still not in conduit. I explained that the lights not being able to operate in sync or independently will be a problem. Fiberstars, the maker of Jazz Lights and the fiber optic system, has a controller that will operate the lights in sync and separate. He said if it could be done he will do it. Fiberstars offers a WPC-3 system that will control their lights and the pumps.
DSC00149Salt system added into pool plumbing. 19 bags of salt added to pool. They hung the controller for the salt system on the small electric conduit that the ozone generator, radio remote, and electrical service is hung on, causing the whole thing to sway in the breeze. Another thing I will have to fix.
Debbie Moles called and said that the fiber optics was done. I told her I had not seen anyone but would look. She called back and said they would be out today. Fiber optic controller, 6004, was placed and turned on. Debbie called and said that they would be back to finish the waterfall fiber optics on Saturday. Wok pot pipes are still blocked with dirt. Wok pot pipes were cut down so they did not stick up out of the pots. The light controller is not an auto sync version as discussed on many occasions. Maybe they will figure it out on their own this time.

Oops

Friday, March 17th, 2006

The electrician repaired the incorrectly wired pumps. He also said that the pool lights were also wired wrong and he fixed that also. He said they also installed the wrong lights in the pool. The color wheels can not be synced together, only one light has the auto sync feature the other light is a standard light. While checking his paperwork he also noted that they had installed Jazz lights, his paper calls for F/X lights. The F/X is a lower output lamp, and on a pool this size why lower output lamp? My contract states Jazz lights. So he needs to call and have a diver come out and replace the lights with the correct ones. They will not be able to operate the lights independently of each other the way they are wired and he doesn’t know if they will be able to sync colors with the fiber optics. According to Fiberstars, http://www.fiberstars.com/, all their lighting should be able to sync with each other assuming you get the right ones.
Jesse with Poco Verde called and said they are not able to seal the wok pots till the fiber optics are done in the pots.
Paddock came by to touch up the chips, stains, and cracks. He looked over the deck and said that he could do the repairs but since the construction was not finished, as he was told, he wanted to wait. He said that he would need to pressure wash and repaint the whole deck. I agreed that we need to wait because they still need to move rock and dirt across the deck.
Prado and Sons Plaster called and will be out on Tuesday to look over the stains in the plaster.
Water level down over 1/2″ since I turned off auto leveler. Pumps did not run all last night because the U/V system was installed.
Called Debbie and left another message.
Debbie Moles called and I gave her the list of things yet to be done. She was under the impression it was complete. She is going to schedule the electricians to figure out the lighting, fiber optic people, Poco Verde to seal the wok pots, Allen Tile to clean up the tile, the deck cleanup, and the plasterer to fix the stains. I am going to give them till the end of March to finish the project, at that time I will have no further contact with them.

Fiber Optics

Wednesday, March 15th, 2006

Fiber optic people were out today. Only thing done was to put wok pot 3’s wire inside conduit. The conduit does not continue where they are going to pour concrete. The wires will be inside the concrete wok pot base. They left the conduit running above ground behind the wok pot bases. The waterfall lighting is lying on top of the waterfall. The PCC2000 automatic cleaning system has failed already. Only half the heads work and even they are pointed in the wrong directions. I believe there is a leak somewhere because the auto water leveler is still running. I shut it off to see how much the water goes down.

Interior

Friday, March 10th, 2006

Interior of pool applied. White 3M Quartz Aggregate applied. Filling pool with water.
Debbie Moles called and said that the pool startup crew would be out on Monday. Fiber optic people would be out on Tuesday or Wednesday. And the deck would be touched up on Thursday or Friday.
Stacey called and said she is having a tech come out Monday to start the equipment. The salt system will be installed in a week or so.

Day 30 of contract – Final Day

Wednesday, February 15th, 2006

Fiber optics in pot #3 replaced. Optics could not be run in conduit. A cut into the concrete at the wok pot base so that the optics line could be rerun into the wok pot. This will place the fiber optics inside the concrete base. If any service is ever needed to the cable, the base would need to be torn up to get to the cable. Debbie Moles called and left a message, she will call again on Thursday.

Pulled

Tuesday, February 14th, 2006

Fiber optics that was broken in pot #3 was pulled from their conduit. I noticed several cracks in the stucco on the face of the waterfall.

Week 46

Thursday, February 9th, 2006

Poco Verde removed wok pots. Fibre Creations came out to place fiber optics. Could not do water sheer due to placement too far into wall. No lip to mount to. Can not cut into the tile due to the light fires down and would be blocked by the tile. Number 3 wok pot fiber optic line broke. Have to re-plumb wok pot fiber optics.

Day 20 of contract

Monday, February 6th, 2006

Front face of waterfall tiled. Side wall is almost complete. Tile sits flush with water sheer. There is no place for the fiber optics to mount.
“The Fiberstars Waterfall Light Bar can be installed
before or after the fall is installed. Note: Installation
of light bar on radius falls before the fall is installed is
recommended. Making sure the bevel on the
light bar track is angling away from the pool wall,
toward the back of the water (see Fig. 2). Simply
attach the light bar to the fall using clear PVC cement.
Be careful not to obstruct the water flow with the light
bar track or PVC cement. Pull fiber through the conduit
at the waterfall back to the illuminator location.
Important! On sheer falls there must be
1 1/8” between the finished tile and the front lip
of the fall. On Rain falls there must be 1 5/8”
clearance for installation of light bar.”

None of these instructions were followed as the water sheer is flush with the tile. These instructions were taken from Fiberstars website, http://www.fiberstars.com.

Week 34

Thursday, November 17th, 2005

Fiber optic people started on adding fiber to waterfall. They cut out the wall from the misplaced conduit to where it was supposed to be, chiseling a line inside the waterfall wall. I spoke with Debbie Moles and told her about the nail holes in the face of the waterfall that Poco Verde was going to repair before painting but did not. I also asked her about the fiber being embedded into the wall. She said that the nail holes would be fixed and that the fiber optics would never be a problem in the wall. She asked me to e-mail her some current pictures of the planter.
E-mail to her.
This is where we are at as of 11-17-05
Picture DSC00075 – front view of planter.
Picture DSC00079 – water sheer with fiber optics cut out?
Picture DSC00080 – I will be happy to buy Poco Verde a level if they need one.

Week 29

Thursday, October 13th, 2005

Debbie Moles called at 11:30am and said Lloyd Wilson was being replaced by Scott Dilliam for now. I told her about the water seeping into the pool from the planter. She said Scott would be out soon to look over the project. She also said Poco Verde was waiting for ‘sheening’ (sheer?) till the top cap could be placed on the wall and that was why the last brick has not been placed.
I met with Lloyd Wilson and Scott Dilliam at 1:15pm. I asked about the top of the wok pedestals being covered with decking and stamped like tile. He said that was the plan. Lloyd said they had to upgrade the waterfall pump to a 2 ½ hp due to the size. He said they were going to float the decking in front of the wok pot pedestals to eliminate the gap in the tile. Poco Verde told him that they could straighten out the wall with stucco. Scott is going to meet with them to discuss what they are going to do about the wall. Heat pump still needs to be connected, pump bases need to be sturdy, fiber optic for wok pots and sheer all still need to be addressed. I was told that the farthest run for fiber optic was about 35 feet and the illuminator could not be moved. He asked about the self closing gates. I reminded him that Paddock was to supply the wrought iron fence and gate. Lloyd was concerned about the 5’ fence requirement since the wok pedestals were close to the common wall. But since we are in Gilbert it should not be a problem.

Planter

Tuesday, September 27th, 2005

Poco Verde continues to work on planter. Fiber optics conduit run, stubbed up on wrong side of wall.

Search

Recent Posts

Archives